For an account that earns simple annual interest,find the interest and the balance of the account.
P = $250, r = 8%,t = 6 years
Any help with this equation would Greatly appreciated.

Answer by checkley71(8403)

I=P(1+R)^T-P
I=250(1+.08)^6-250
I=250(1.58687)-250
I=396.72-250
I=$146.72 IS THE INTEREST.
250+146.72=$396.72 IS THE ACCOUNT BALANCE.